OpenStudy (anonymous):

factor the polynomial x^3+4x^2+16x+64

OpenStudy (anonymous):

$$x^3+4x^2+16x+64=x^2(x+4)+16(x+4)=(x+4)(x^2+16)= $$ $$=(x+4)(x+2i)(x-2i) $$

OpenStudy (anonymous):

Oooooo! There is a mistake=(

OpenStudy (anonymous):

The last line $$ =(x+4)(x+4i)(x-4i)$$